A) Create an example of a math problem that says to, “Factor fully” and that includes
Common Factoring of any polynomial.

B) Explain/Teach in detail what you had to do to arrive at your solution. Show all steps! l
1
Expert's answer
2020-12-15T03:34:32-0500

"\displaystyle

 \textsf{Factorize}\,\, x^3 - 6x^2 + 11x - 6 = 0\\

\textsf{Let}\,\, f(x) = x^3 - 6x^2 + 11x - 6 = 0\\

\textsf{Substitute}\,\, x = 0, \pm1, \pm2,\cdots\\

f(0) = -6,\,\, f(1) = 1^3 - 6(1^2) + 11(1) - 6 = 1 - 6 + 11 - 6 = 0\\

\textsf{Since}\,\, f(1) = 0,\\ \textsf{By Factor theorem}\,\, x - 1\,\, \textsf{is a factor of the polynomial}\\

\textsf{Divide the polynomial}\\
\textsf{by the factor}\\

x^2 - 5x + 6\\
x - 1|x^3 - 6x^2 + 11x - 6 = 0\\
-(x^3 - x^2)\\
= -5x^2 + 11x - 6\\
-(-5x^2 + 5x)\\
= 6x - 6\\
-(6x - 6)\\
= 0\\


\textsf{Factorize the quotient}\\

x^2 - 5x + 6 = (x - 3)(x - 2)\\

f(x) = \textsf{divisor}\times \textsf{quotient}\\

\textsf{Therefore,}\,\,x^3 - 6x^2 + 11x - 6 = (x - 1)(x - 2)(x - 3)"